### v1

# Self-consistency error correction for accurate machine learning potentials from variational Monte Carlo

Abstract
Variational Monte Carlo (VMC) can be used to train accurate machine learning interatomic potentials (MLIPs), enabling molecular dynamics (MD) simulations of complex materials on time scales and for system sizes previously unattainable. VMC training sets are often based on partially optimized wave functions (WFs) to circumvent expensive energy optimizations of the whole set of WF parameters. However, frozen variational parameters lead to VMC forces and pressures not consistent with the underlying potential energy surface, a bias called the self-consistency error (SCE). Here, we demonstrate how the SCE can spoil the accuracy of MLIPs trained on these data, taking high-pressure hydrogen as test case. We then apply a recently introduced SCE correction [Phys. Rev. B 109, 205151 (2024)] to generate unbiased VMC training sets based on a Jastrow-correlated single determinant WF with frozen Kohn-Sham orbitals. The MLIPs generated within this framework are significantly improved and can approach in quality those trained on datasets built with fully optimized WFs. Our conclusions are further supported by MD simulations, which show how MLIPs trained on SCE-corrected datasets systematically yield more reliable physical observables. Our framework opens the possibility of constructing extended high-quality training sets with VMC.
1 Introduction
Since their introduction 1, 2, machine learning interatomic potentials (MLIPs) have revolutionized the field of molecular dynamics (MD) 3, 4, 5, 6, 7, 8, 9, 10, 11, 12. These models can learn the potential energy surface (PES) of a given system starting from a representative set of configurations, called the training set. Each element of this set is labeled by its physical properties (e.g., energy, ionic forces, pressure, etc.), usually computed using ab initio electronic structure methods. After it is trained, an MLIP is capable of reproducing the observables of the system on new configurations with the accuracy of the original method but with a computational cost of several orders of magnitude smaller 2. This has opened the possibility of running MD simulations at ab initio accuracy with sizes and lengths that were previously out of reach 13. By far, the most common method used for generating MLIPs datasets is density functional theory (DFT) 14, which can easily produce $\sim 10^{4}$-$10^{5}$ training points with modest computational resources, usually sufficient to train data-heavy machine learning frameworks 7, 15. For several systems, however, DFT fails to capture the strong effect of electronic correlations or may provide results that strongly depend on the choice of the exchange-correlation functional. Quantum Monte Carlo (QMC) methods are a series of stochastic algorithms that provide state-of-the-art results for a large class of systems, from molecules to solids 16, 17. Although more accurate, QMC approaches have a much larger computational cost than DFT. In the context of MLIPs, this implies that the number of training configurations that can be generated within a given amount of resources is generally smaller than the one attainable with mean field methods. Recently proposed solutions to this problem make use of data-efficient schemes in combination with hierarchical machine learning (or $\Delta$-learning 18) to construct QMC-based MLIPs 19, 20. However, within the most popular QMC approach, i.e., variational Monte Carlo (VMC), another issue prevents a straightforward application of MLIPs, presented in what follows.
The efficiency of VMC is largely due to the complexity of the wave function (WF) employed in the calculations to describe the electronic part. A possible choice for the WF $\Psi$ is
$\Psi=\exp\left(J\right)\times\Phi_{\textrm{AS}},$
where $\exp\left(J\right)$ is a bosonic term (i.e., symmetric under particle exchange) called the Jastrow factor, and $\Phi_{\textrm{AS}}$ is an antisymmetric (AS) part. Both terms contain variational parameters, whose number can lead to flexible WFs. However, there is a practical difficulty in optimizing a large number of variables in a stochastic setting, particularly those belonging to the AS part. Thus, only the Jastrow parameters are usually optimized, and $\Phi_{\textrm{AS}}\equiv\Phi_{\textrm{SD}}$ is often taken as a single Slater determinant (SD) obtained from a DFT calculation 17. This gives rise to the frozen-orbital Jastrow Slater determinant (JSD) wave function, an appealing candidate for VMC-based machine learning applications, since it provides an excellent tradeoff between accuracy and computational cost. Indeed, the Jastrow optimization is much less demanding than the full optimization also involving the determinant. Nevertheless, the use of the frozen Slater determinant introduces a bias in the VMC forces and pressures computed with this WF. This bias has been dubbed as the self-consistency error  21, 22, and we are going to show that it is particularly relevant in the context of VMC-based MLIPs. Even though machine learning potentials are, by definition, consistent, a biased training set can nonetheless spoil the final accuracy of the model. A solution would be to use only the energies for the training; this, however, usually increases by orders of magnitude the number of configurations necessary to reach a given accuracy 23, 24.
In Ref. 22, a suitable correction was developed to remove the self-consistency error, adding a relatively small computational overhead to the calculation. The aim of this work is to show the actual impact of the self-consistency error on an MLIP in a working case scenario, and to demonstrate how its correction can be applied to generate unbiased VMC datasets with a JSD WF for machine-learning applications. As test case, we consider high-pressure hydrogen, a widely studied system with applications ranging from planet modeling 25 to inertial confinement fusion 26, 27, 28, and often used to benchmark numerical methods 29. In particular, we focus on density and temperatures relevant for the Hugoniot curve, i.e., the set of possible states reachable with a shock wave 30. The deuterium Hugoniot was recently computed in Ref. 20 using MLIPs trained on accurate QMC data. There, the self-consistency error was avoided by relaxing the full set of variational parameters in the WF, i.e., including both the Jastrow factor and the antisymmetric part. The resulting optimization improves the consistency of the forces and pressures and also changes the underlying PES, at a computational cost about 10 times larger than the standard frozen-orbital JSD optimization.
Here, we explore a different path, by showing that a very similar accuracy can be obtained by training models on a frozen-orbital JSD dataset supplemented by the self-consistency error correction. In particular, we compute VMC forces and pressures on the same configurations using a JSD WF, with and without corrections, and use them to train the different models. The performances of these MLIPs are compared with the one of an MLIP trained on the fully optimized dataset of Ref. 20, which we take as the reference. In this way, we show the relevance of the self-consistency error in the MLIP generation and the need for its correction, by studying the impact of both the correction and the full optimization of the WF on thermodynamic equilibrium quantities such as pressure and radial distribution functions. This work demonstrates the importance of obtaining unbiased QMC datasets for MLIPs at the VMC level using cheap-to-optimize JSD-type WFs, including also ionic forces and stress components.
2 Methods
We briefly present the main methods used in this work. In Sec. 2.1 we introduce general aspects of MLIPs, focusing on one specific framework, i.e., the kernel ridge regression, and on the loss function used for model training. In Sec.2.2, we review the forces and pressure correction developed in Ref. 22 to solve the self-consistency error in VMC and describe its implementation in our WF representation.
2.1 Generalities of MLIPs
The basic goal of any MLIPs is to predict the energy of a given $N$-atom configuration described by its ionic positions, i.e., $E=E\left(\mathbf{R}_{1},\cdots,\mathbf{R}_{N}\right)$. We now suppose that $E$ can be expressed as a sum of atomic contributions 2, each depending on the relative coordinates of all the other atoms with respect to the $i$-th one:
$E=\sum_{i=1}^{N}e(\mathcal{R}_{i})\qquad\text{where}\quad\mathcal{R}_{i}=\left(\mathbf{R}_{i1},\dots,\mathbf{R}_{iN}\right)\quad\text{with}\quad\mathbf{R}_{ij}\equiv\mathbf{R}_{j}-\mathbf{R}_{i}.$ (1)
Here, we refer to $\mathcal{R}_{i}$ as the local environment around the $i$-th atom and to $e\left(\mathcal{R}_{i}\right)$ as the local atomic energy, and suppose the minimum image convention holds when periodic boundary conditions are applied. Often, the function $e\left(\mathcal{R}_{i}\right)$ in Eq.(1) is considered local, i.e., only dependent on the atoms closer than a certain cutoff radius $r_{c}$. Different types of MLIPs may be distinguished by their internal representation of the local atomic energy. Many techniques have been successfully applied over the years, including deep neural networks 2, 4, 11, Gaussian approximation potentials 31, 32, kernel ridge regression 33, 19, graph neural networks 34, 35, and more. Independently of the specific architecture, $e\left(\mathcal{R}_{i}\right)$ usually depends on a set of model parameters $\mathbf{p}$ which have to be optimized:
$e\left(\mathcal{R}_{i}\right)\equiv e\left(\mathcal{R}_{i};\mathbf{p}\right).$ (2)
Most MLIPs implement a functional form that already satisfies some of the symmetries required by physical constraints, such as permutational and rotational invariance. To determine the model parameters $\mathbf{p}$, we consider a set of $N_{\textrm{train}}$ configurations, i.e., the training set, for which we compute the desired physical quantities using a target numerical method. In particular, we indicate with $\mathcal{R}^{\mu}_{i}\equiv(\mathbf{R}^{\mu}_{1}-\mathbf{R}^{\mu}_{i},\dots,\mathbf{R}^{\mu}_{N_{\mu}}-\mathbf{R}^{\mu}_{i})$ the $i$-th atomic environment belonging to the $\mu$-th configuration of the set, with $\mu=1,\dots,N_{\textrm{train}}$, having $N_{\mu}$ atoms and volume $V_{\mu}$ (for simplicity we will consider the supercell to be cubic). Moreover let $E^{\mu}_{\textrm{ref}}$, $\mathbf{F}_{j,\textrm{ref}}^{\mu}$ and $P^{\mu}_{\textrm{ref}}$ be the values of the total energy, the force acting on the $j$-th atom, and the (isotropic) virial pressure of the configuration. The model parameters $\mathbf{p}$ are then optimized by minimizing a loss function such as
$\mathcal{L}(\mathbf{p})$ $=W_{E}\frac{1}{N_{\textrm{train}}}\sum_{\mu=1}^{N_{\textrm{train}}}\left[\frac{1}{N_{\mu}}\sum_{i=1}^{N_{\mu}}e\left(\mathcal{R}_{i}^{\mu};\mathbf{p}\right)-\frac{1}{N_{\mu}}E^{\mu}_{\textrm{ref}}\right]^{2}$
$+W_{F}\frac{1}{N_{\textrm{train}}}\sum_{\mu=1}^{N_{\textrm{train}}}\frac{1}{3N_{\mu}}\sum_{j=1}^{N_{\mu}}\sum_{\alpha=x,y,z}\left[-\frac{\partial}{\partial R^{\mu}_{j,\alpha}}\sum_{i=1}^{N_{\mu}}e\left(\mathcal{R}_{i}^{\mu};\mathbf{p}\right)-F^{\mu}_{j,\alpha,\textrm{ref}}\right]^{2}$
$+W_{P}\frac{1}{N_{\textrm{train}}}\sum_{\mu=1}^{N_{\textrm{train}}}\left[-\frac{\partial}{\partial V_{\mu}}\sum_{i=1}^{N_{\mu}}e\left(\mathcal{R}_{i}^{\mu};\mathbf{p}\right)-P^{\mu}_{\textrm{ref}}\right]^{2}.$ (3)
The three terms in Eq.(3) are the mean squared error (MSE) of the energy per atom, ionic forces, and virial isotropic pressure, respectively, and $W_{E}$, $W_{F}$, and $W_{P}$ are tunable weights multiplying the different MSEs. Other functional forms of the loss function are possible, including additional physical quantities (e.g., charges, all $6$ independent components of the stress, etc). In neural networks, $\mathcal{L}(\mathbf{p})$ can be minimized using stochastic optimization methods that compute the gradients using only a small batch of training configurations. In this paper, we primarily use MLIPs based on the kernel ridge regression (KRR) method. Within KRR, the local atomic energy is expressed as
$e\left(\mathcal{R};\mathbf{p}\right)=\sum_{\nu=1}^{N_{\textrm{env}}}p_{\nu}K\left(\mathcal{R},\mathcal{R}_{\nu}\right)$ (4)
where $\mathcal{R}_{\nu}$, $\nu=1,\dots,N_{\textrm{env}}$ belong to a subset of all the local environments in the training set, and $K(\mathcal{R},\mathcal{R_{\nu}})$ is the normalized kernel between $\mathcal{R}$ and $\mathcal{R}_{\nu}$. Following Refs 19 and 20, we adopt a kernel based on a variant of the Smooth Overlap of Atomic Positions (SOAP) descriptor 36, 37(see App. A.1), and we select the $N_{\textrm{env}}$ local environments using the farthest point sampling method 19, according to the distance introduced by the kernel $K$. Notice how plugging Eq.(4) into the loss function (3) and taking the derivatives with respect to the variational coefficients turns the minimization problem into a linear system, which is then solved by the conjugate gradient method. We conclude this Section by stressing the fact that, by definition, any MLIPs will give energy derivatives (e.g., forces and virial stress components) that are consistent with their underlying potential energy surface. Any effect of potential inconsistency in the dataset, such as the ones studied in this work, will manifest in the final performance of the models and its dependence on the relative training weights $W_{E}$, $W_{F}$, and $W_{P}$. This is demonstrated in Sec.3.2, where we compare the quality of MLIPs trained on different datasets as a function of $W_{E}$, $W_{F}$, and $W_{P}$.
2.2 Self-consistency error in VMC and its correction
Within VMC, the force acting on the $i$-th atom is computed using
$\mathbf{F}^{\textrm{VMC}}_{i}=-\left\langle\frac{\partial}{\partial\mathbf{R}_{i}}E_{L}\right\rangle-2\left\langle(E_{L}-E)\frac{\partial}{\partial\mathbf{R}_{i}}\log\Psi\right\rangle,$ (5)
[…]
$\mathbf{F}^{\textrm{corr}}_{i}=-\sum_{k=1}^{N_{SD}}\frac{\partial E}{\partial\alpha^{\textrm{SD}}_{k}}\frac{d\alpha^{\textrm{SD}}_{k}}{d\mathbf{R}_{i}}$ (6)
has to be considered, containing the derivatives of the $N_{\textrm{SD}}$ variational coefficients $\alpha_{k}^{\textrm{SD}}$ included in the Slater determinant.
Discarding the term in Eq.(6) causes an inconsistency between the forces and the underlying PES computed with a JSD WF, meaning that $\mathbf{F}^{\textrm{VMC}}_{i}\neq-\frac{\mathop{dE}}{\mathop{d\mathbf{R}_{i}}}$. This is called the self-consistency error (SCE) 21, 22. An analogous discussion applies to other energy derivatives, such as the virial pressure, that contain a similar bias and for which we can derive similar corrections.
In our applications, following the implementation of the TurboRVB package 45 (here used for all the QMC calculations), we consider a representation of the SD part in terms of the antisymmetrized geminal power (AGP)
$\Phi_{\textrm{AGP}}=\mathcal{A}\left(g(x_{1},x_{2})\cdots g(x_{N-1},x_{N})\right),$ (7)
where $x_{i}=(\mathbf{r}_{i},\sigma_{i})$ indicates the collective spatial and spin coordinates of the $i$-th electron, with $i=1,\dots,N$, $\mathcal{A}$ is the antisymmetrizer operator and $g$ is a pairing function. In Eq. (7) we also suppose $N$ to be even for simplicity. In the spin unpolarized case, the pairing function can be expressed as the product of a spatial part $f$ and a spin singlet, namely
$g(x_{i},x_{j})=f(\mathbf{r}_{i},\mathbf{r}_{j})\frac{|\uparrow\downarrow\rangle-|\downarrow\uparrow\rangle}{\sqrt{2}}.$ (8)
In TurboRVB the spatial part of the pairing function is written in terms of a localized basis of $L$ atomic orbitals (AOs) $\phi_{k}$
$f(\mathbf{r}_{i},\mathbf{r}_{j})=\sum_{k=1}^{L}\sum_{k^{\prime}=1}^{L}\lambda_{k,k^{\prime}}\phi_{k}^{\theta}(\mathbf{r}_{i})\phi_{k^{\prime}}^{\theta^{\prime}}(\mathbf{r}_{j}).$ (9)
In our case, the AOs are taken as Gaussian type orbitals (GTOs). For periodic systems, such as the ones considered in this work, the localized basis is appropriately generalized to satisfy periodic or twisted boundary conditions 46, depending on the phase $\theta$ ($\theta^{\prime}$) for spin-up (spin-down) electrons. Here, $\theta=-\theta^{\prime}$, such that the global phase factor is canceled, making the function $f$ invariant under a global translation. Fixing this gauge leads to a numerically stable evaluation of the $d\lambda_{k,k^{\prime}}/d\mathbf{R}$ derivatives implied by Eq. 6.
An alternative expression for the function $f$ can also be obtained using molecular orbitals (MOs) $\Phi_{n}(\mathbf{r})=\sum_{k}c_{n,k}\phi_{k}(\mathbf{r})$. In particular, when the number $M$ of MOs is equal to $N/2$, the resulting AGP is exactly equivalent to a single SD, i.e., $\Phi_{\textrm{AGP}}\equiv\Phi_{\textrm{SD}}$. Within TurboRVB, the MOs (and the corresponding $\lambda_{k,k^{\prime}}$ matrix) are initialized by a local density approximation 47 (LDA) DFT calculation. This is performed in a basis consisting of the aforementioned AOs with an additional one-body Jastrow factor to automatically take into account Kato cusp conditions 48, 45. Because of this, the electronic integrals entering into the Kohn-Sham equations are evaluated on a real space grid with a suitably chosen lattice space. Finally, we also mention that the basis is regularized by removing any eigenvector of the overlap matrix $S_{k,k^{\prime}}=\langle\phi_{k}|\phi_{k^{\prime}}\rangle$ with a corresponding eigenvalue ($s_{i}$) satisfying the condition $s_{i}/s_{\rm max}<\epsilon_{\textrm{cut}}$ 49, where $s_{\rm max}$ is the maximum eigenvalue and $\epsilon_{\textrm{cut}}$ is set to $10^{-7}$ in this study.
Finally, the expression in Eq. (6) is stochastically evaluated in VMC according to
[…]
where $\mathcal{O}_{k,k^{\prime}}=\frac{\partial\log\Psi}{\partial\lambda_{k,k^{\prime}}}$ and $\overline{\mathcal{O}}_{k,k^{\prime}}=\langle\mathcal{O}_{k,k^{\prime}}\rangle$. In the previous expression, the logarithmic derivatives in $\mathcal{O}_{k,k^{\prime}}$ are efficiently computed using the adjoint algorithmic differentiation 42, while the derivatives of the parameters, $\frac{d\lambda_{k,k^{\prime}}}{d\mathbf{R}_{i}}$, are evaluated within DFT-LDA using the finite differences method (FDM), as described in Ref. 22. For the virial isotropic pressure, an analogous derivation yields the following expression for the correction term:
[…]
3 Results
Henceforth, we analyze how the SCE affects the training of MLIPs and discuss how the correction described in Sec. 2.2 can effectively be implemented to improve the accuracy of the resulting models. With this aim, we consider a dataset made of pristine hydrogen configurations for which we computed VMC energy, forces, and pressure using different types of WF and with/without bias correction. The dataset is the same as the one used in Ref. 20 and comprises $558$ configurations of $128$ atoms each. These configurations were extracted from DFT-driven MD simulations at temperatures between $4000$ K and $20000$ K and Wigner-Seitz radii $r_{s}$ between $1.80$ and $2.12$. This range of thermodynamic conditions corresponds to the expected location of the Hugoniot curve of deuterium, defined as the set of states that can be reached by producing a shock wave in a sample. For these configurations, we fully optimized the wave function, going beyond the frozen-orbital JSD ansatz. As previously mentioned, we take this dataset as the reference one. In Sec. 3.1, we study the SCE for the same set of configurations as the reference dataset, and show the effect of both forces and pressure corrections. In Sec. 3.2 we directly compare the accuracy of MLIPs trained on a biased dataset (i.e. affected by the SCE) and a corrected one, respectively. We also compare both models with the reference dataset used in Ref. 20. Finally, in Sec. 3.3, we analyze the results of molecular dynamics simulations driven by the different MLIPs.
3.1 Correcting the SCE bias in the Hugoniot dataset
To assess the magnitude of the SCE in the Hugoniot dataset, we use a JSD WF with a basis set of [4s2p1d] and [2s2p1d] GTOs for the antisymmetric part and Jastrow factor, respectively. The same primitive basis set was used in Ref. 20. We initialize the SD by running an LDA-DFT calculation with twisted boundary conditions at the Baldereschi point, i.e, at $\mathbf{k}=\left(\frac{1}{4},\frac{1}{4},\frac{1}{4}\right)$ in crystal units. The Jastrow part is optimized using the stochastic reconfiguration method 50.
We assess the consistency of the JSD forces and pressures by calculating the forces and pressure on a few selected configurations with two separate approaches: (i) evaluating Eq.(5) (along with the analogous expression for pressure) and (ii) using the FDM by fitting the PES of the system and evaluating the derivative numerically. As previously discussed, the SCE manifests itself as a difference between these two values. The results are shown in Fig. 1 for one of the force components and the virial pressure.
We observe a significant bias for both quantities, with a discrepancy as large as $\sim 10\%$ of their values. This clearly highlights the importance of curing the SCE for the biased dataset. For this reason, we applied the corrections of Eqs. (10) and (11) to forces and pressures, respectively. For forces, we computed the numerical derivatives of the parameters $\frac{d\lambda_{k,k^{\prime}}}{d\mathbf{R}}$ with the FDM from DFT-LDA, the theory used to obtain $\lambda_{k,k^{\prime}}$, with a displacement $\Delta R=\pm 0.003$. The derivatives of $\lambda_{k,k^{\prime}}$ with respect to the volume, entering in the pressure correction, were obtained with the FDM using relative volume variations of ±$\pm 0.03\%$. In Fig. 1 we observe that the corrected forces and pressure are perfectly compatible with their values directly computed through the fit of the VMC PES. Fig. 2 shows the values of both the biased and corrected force components (pressures) versus the forces (pressures) estimated from the PES, for several configurations. Notice how the pressure correction acts almost like a rigid shift of $\sim 1$ GPa. The root mean squared error (RMSE) of biased and corrected quantities is also shown, further demonstrating the effectiveness of the correction in removing the SCE.
Sometimes, the added correction significantly increases the error bar $\sigma_{F}$ of the corrected forces. We then disregard the corrections that lead to forces with $0.0153\sigma_{F}>$ 0.015 Ha/Bohr. For pressures, the threshold value of $1.5\times 10^{-5}$ is applied to $3\sigma_{p}$. See App. A.2 for further details. In principle, the FDM used to estimate the parameter derivatives $\frac{d\lambda_{k,k^{\prime}}}{d\mathbf{R}}$ can be replaced with more robust and faster approaches (e.g., using linear response theory 51). We leave this possibility for future work.
Figure 1: (a) Comparison among the biased pressure evaluated with an expression equivalent to Eq. (5) (red diamonds), the corrected pressure obtained by applying Eq. (11) (violet square), and the pressure obtained by fitting the PES (dashed green line and green diamond) for a selected configuration in the Hugoniot dataset20. The PES of the system is also shown (green dots and dash-dotted line). (b) Comparison among the biased force evaluated with Eq. (5) (red diamonds), the corrected force obtained by applying Eq. (10) (violet square), and the force calculated by fitting the PES (dashed green line and green diamond) for the $z$-component of the second hydrogen atom belonging to the same configuration as in (a). The PES of the system along the atomic displacement is also shown (green dots and dash-dotted line).
Figure 2: (a) Value of the biased (red markers) and corrected (violet markers) pressure for $6$ different $128$-atom configurations as a function of the numerical pressure estimated from the derivative of the PES with respect to the volume. (b) Values of the biased (red markers) and corrected (violet markers) force components as a function of the numerical force estimated by the PES for one of the configurations belonging to the Hugoniot dataset20. The dashed line is a reference for perfect consistency.
3.2 Comparison between VMC datasets
To illustrate how using an SCE-corrected JSD dataset affects the MLIPs generation, we compare the performances of MLIPs trained on three different datasets: (i) the aforementioned "biased dataset" obtained with a JSD WF and with forces and pressure computed with the standard expressions (Eq. 5), (ii) a "corrected dataset" where the corrections of Eqs. (10) and (11) are applied as described in Sec. 2.2 and (iii) a "reference dataset", i.e., the one introduced in Ref. 20. In particular, the latter was obtained from an improved Jastrow-correlated AGP WF, where the Jastrow and antisymmetric part were both optimized, thanks to a combination of geminal embedded orbitals52 to reduce the number of variational parameters and a restricted optimization based on the locality of the AGP $\lambda_{k,k^{\prime}}$ matrix. The $\lambda_{k,k^{\prime}}$ coefficients were optimized for those basis set elements $k$ and $k^{\prime}$ centered on atoms closer than a cutoff distance of 4 Bohr radii20. No significant lowering of the variational energy was found beyond this threshold, suggesting that the dependence of the energy on the remaining parameters can be disregarded. Thanks to this approach, the SCE on this dataset was effectively removed and we thus decided to take the models trained on it as the reference. Notice how the explicit optimization of the whole WF, as the one performed in Ref. 20, is a procedure that is about an order of magnitude more expensive than the partial optimization of the frozen-orbital JSD WF, and requires advanced minimization techniques 53, 54.
For all the MLIPs, a KRR model is trained on the difference between VMC quantities and DFT ones, the latter using the Perdew-Burke-Ernzerhof (PBE) exchange-correlation functional 55, following the $\Delta$-learning approach 18, 19, 20. In this way, a much higher accuracy is achieved with relatively small datasets such as the ones used here. The values of the hyperparameters chosen for these models are reported in App. A.1. Among the $558$ configurations of the datasets, we select $58$ configurations for testing. We investigate the effect of varying the training weights in the loss function (Eq. (3)) on the RMSE computed for energy, forces, and pressure of the test set for each dataset.
Since our goal here is to compare MLIPs trained on different datasets (i.e., different training values for energy $E$, forces $\mathbf{F}$ and pressure $P$), the RMSE may not be the best metric to assess the relative accuracy of one model with respect to the others. For this reason, we define the relative improvement of the model for each system property ($X=E,\mathbf{F},P$) as:
$\Delta_{X}=\frac{\mathrm{RMSE}(X_{\textrm{test}}-X_{\textrm{pred}})-\mathrm{RMSE}(X_{\textrm{test}}-X_{\textrm{dummy}})}{\mathrm{RMSE}(X_{\textrm{test}}-X_{\textrm{dummy}})},$ (12)
where $X_{\textrm{pred}}$ are the model predictions, $X_{\textrm{test}}$ the test set values, $E_{\textrm{dummy}}=\frac{1}{N_{test}}\sum_{i\in test}E_{i}$, $\mathbf{F}_{\textrm{dummy}}=0$, and $P_{\textrm{dummy}}=0$. In other words, $\Delta_{X}$ measures the relative improvement of the model prediction on the quantity $X$ with respect to a "dummy" model, corresponding to a perfectly flat PES equal to the average energy on the test set for all configurations. Notice how $\Delta_{X}=-1$ indicates that the model has learned exactly the quantity $X$, while a value $\Delta_{X}\geq 0$ means that the model has an error equal to, or larger than, a flat model. In the present work, the test quantities refer to the difference between the VMC and the DFT-PBE baseline, while $E_{\textrm{dummy}}$ is the average difference on the test set.
Figs. 3 and 4 show the results obtained for $\Delta_{E}$, $\Delta_{F}$, and $\Delta_{P}$ using MLIPs trained with different values of the training weights $W_{E}$, $W_{F}$, and $W_{P}$. We first analyze what happens by progressively increasing the weight on the forces for the different models without explicitly training on pressures (i.e., $W_{P}=0$). The force bias in the biased dataset is evident from the larger $\Delta_{F}$ of the corresponding model compared to that of both the corrected and reference datasets (see Fig. 3). Notice how, for small values of $W_{F}/W_{E}$, $\Delta_{F}$ is positive in the biased case, which in the $\Delta$-learning framework means that the correction on forces is detrimental. As the weight on forces is increased, all models reach a plateau in $\Delta_{F}$, while simultaneously losing some accuracy on the other quantities. For the energy, this loss of accuracy is significantly larger in the models trained on the biased dataset, further demonstrating the effect of the SCE.
We observe very similar performances between the corrected and reference models. For pressure, the models including the SCE correction show the best performances for almost all the values of $W_{F}$. Since the loss does not include the pressure in this case, this is another indication of consistency. In Fig. 4 we plot the results obtained by varying the weight on pressure, for two values of $W_{F}/W_{E}$, namely $W_{F}/W_{E}=3/128$ and $W_{F}/W_{E}=15/128$. Also in this case, a better accuracy is generally achieved for the models trained on the corrected and reference datasets.
Finally, we mention how these conclusions exclusively depend on the quality of the datasets themself, i.e., their internal consistency, and not on the MLIPs specific architecture. To demonstrate this feature, in App. A.3, we discuss the relative improvement $\Delta_{X}$ for models obtained with the MACE 35 framework, which show the same behavior as the one found for the KRR MLIPs.
Figure 3: Relative RMSE variation $\Delta_{X}$ (Eq. 12) for energy, forces, and pressure, as a function of the force/energy weight ratio in the loss function of Eq. 3 (expressed in atomic units) for MLIPs trained on the biased, corrected and reference datasets, respectively. The weight on the pressure is set to zero for all models.
Figure 4: $\Delta_{X}$ (Eq. 12) for energy, forces, and pressure, as a function of the pressure/energy weight ratio in the loss function of Eq. 3 used for training (expressed in atomic units), for the different models. Solid lines correspond to $W_{E}=1$ and $W_{F}=3/128$, while dashed lines correspond to $W_{E}=1$ and $W_{F}=15/128$. The arrows highlight the "best" MLIPs for each of the three datasets, which are used for the MD simulations described in Sec. 3.3.
3.3 MD results comparison
To conclude our analysis, we run a set of MD simulations using MLIPs trained on the three different datasets. For each model, we select the relative training weights $W_{F}/W_{E}$ and $W_{P}/W_{E}$ to have a good tradeoff between the RMSE values on energy, forces, and pressure. In particular, for the MLIPs trained on the reference and corrected weight, we choose the last value of $W_{F}/W_{E}$ before the significant increase in $\Delta_{E}$ (i.e., the fourth smallest value in Fig. 3), in order to increase as much as possible the accuracy on forces without spoiling too much the energy. A similar reasoning is used to determine the optimal value of $W_{P}/W_{E}$ (see Fig. 4). For the MLIP trained on the biased dataset we choose a larger value of $W_{F}/W_{E}$, which is necessary to reach a reasonable accuracy on forces. This comes at the price of sacrificing the accuracy on the energy, a direct consequence of the SCE. The weights used for each MLIP are summarized in Tab. 1.
Dataset $W_{F}/W_{E}$ $W_{P}/W_{E}$
Biased $15/128$ $10^{6}/128^{2}$
Corrected $3/128$ $10^{6}/128^{2}$
Reference $3/128$ $10^{6}/128^{2}$
Table 1: Training weights for the MLIPs used in the MD simulations, for each dataset. For both the forces and pressure weights, we report the relative value with respect to $W_{E}$. All values are in atomic units.
We run MD simulations at three different density and temperature conditions close to the position of the deuterium Hugoniot curve given by the VMC-MLIP of Ref. 20. In particular, we considered thermodynamic conditions which span both the molecular liquid and the atomic one. Because of the $\Delta$-learning approach adopted, the resulting efficiency of these calculations is the same as a DFT Born-Oppeheimer MD simulation. Indeed, the energy, ionic forces and pressure of the system are calculated at each step with DFT, and the corresponding corrections are then added using the MLIP. Notice how the DFT baseline can be, in principle, replaced by a faster potential (e.g., using an MLIP trained on DFT data 56), improving the overall efficiency. Further details regarding the computational setup of the MD simulations are given in App. A.4.
A comparison of the equilibrium pressure during the dynamics for three different temperatures is reported in Tab. 2. Notice that, in principle, the reference dataset and the corrected one can give different results. Indeed, the optimization of the antisymmetric part of the WF not only improves the consistency of forces and pressure, but can also modify the PES of the system. Nevertheless, if compared with the biased model, the MLIP trained on the corrected data gives results that are consistently much closer to the reference ones. This suggests that the main discrepancy between the biased and reference models comes from the SCE, and that the SCE correction improves considerably the physical description of the system. We observe a maximum discrepancy of $\sim 2$ GPa at $r_{s}=2.02$ and $T=6000$ K between the biased and corrected results, while the corrected and reference ones are in statistical accordance within a joint error bar of 0.2 GPa at the same thermodynamic conditions. Remarkably, this state is very close to the Hugoniot compressibility peak, i.e., the maximum in the density reached by the shocked state along the Hugoniot curve. This thermodynamic point is very sensitive to changes in the underlying equation of state, and a pressure shift as the one observed here between the corrected and biased models has a significant impact in the final position of the Hugoniot 57. Similar conclusions can be reached by looking at the radial distribution function $g(r)$ (Fig. 5). Analogously to the pressure analysis, the $g(r)$ of the corrected model is compatible with the one given by the reference one. Also for this quantity the maximum discrepancy between the corrected and biased models is observed close to the compressibility peak at $T=6000$ K, which corresponds to the onset of the the molecular-to-atomic transition along the Hugoniot curve. These results further highlight how SCE forces and pressure corrections significantly improve the physical outcome of the resulting MLIPs.
$P_{\textrm{bias}}$ (GPa) $P_{\textrm{corr}}$ (GPa) $P_{\textrm{ref}}$ (GPa)
$r_{s}=2.16$, $T=4000$ K $17.1(1)$ $17.4(1)$ $17.4(1)$
$r_{s}=2.02$, $T=6000$ K $26.8(1)$ $28.5(1)$ $28.7(1)$
$r_{s}=1.92$, $T=8000$ K $38.3(1)$ $39.8(1)$ $39.9(1)$
Table 2: Average pressure from MD simulations obtained with different MLIPs trained on the biased, corrected, and reference datasets, respectively. All models employ the $\Delta$-learning technique with a DFT-PBE baseline.
Figure 5: Radial distribution function $g(r)$ for densities and temperatures close to the Hugoniot curve, obtained with MLIPs trained on different datasets. The difference $\Delta g(r)$ with respect to the model trained on the reference dataset is also shown for both the "biased" and "corrected" models. The shaded area indicates the statistical uncertainty in $\Delta g(r)$.
4 Conclusions
Machine learning models have become a key tool to reach system sizes and simulation times not achievable with first principle methods, while showing a similar accuracy. In particular, MLIPs trained on data generated by advanced quantum-mechanical algorithms, such as quantum Monte Carlo (QMC), are essential to improve the description and our physical understanding of phenomena for which standard mean-field methods fail. In this work, we provided evidence on how VMC, in combination with a cheap-to-optimize JSD WF and a force/pressure correction, can be used to construct reliable QMC-based training sets. Our results can be summarized in the following three points.
First, we clearly demonstrated how the SCE affecting derivative quantities (i.e., forces and pressures) computed with a partially optimized WF, such as the frozen-orbital JSD, can heavily spoil the accuracy of MLIPs trained on those data. This is manifested by a large dependence of the models’ accuracy on the relative weights in the loss function.
Second, we showed how an easy-to-evaluate bias correction, previously developed in Ref. 22, can be applied to remove the SCE from the training set, and improve the accuracy of the resulting MLIPs.
Finally, we directly compared equilibrium thermodynamic quantities yielded by MD simulations driven by different MLIPs for the deuterium Hugoniot curve, taken as test system. We observed how the main discrepancy between MLIPs obtained from frozen-orbital and fully optimized WF datasets does not necessarily come from a different PES description of the dataset but from the intrinsic SCE affecting forces and pressures computed with partially optimized WFs. Indeed, in our working case provided by high-pressure hydrogen, the MLIPs trained on the corrected dataset give results in statistical agreement with those yielded by a reference model trained on VMC data with a fully optimized WF.
Our results reveal the importance of appropriately curing forces and pressures when they are computed with partially optimized WF in QMC and subsequently used to train MLIPs. The MLIPs quality degradation due to the SCE plaguing these VMC datasets is an effect that has been overlooked in previous applications. This could be even more relevant for heavier elements, since the SCE usually grows with the number of valence electrons 21. The availability of an SCE-correction framework used in combination with other approaches, such as $\Delta$-learning, to further increase the precision of the final models, opens the way to the efficient generation of accurate MLIP based on QMC electronic structure calculations.
Acknowledgments
K.N. is grateful for computational resources from the Numerical Materials Simulator at National Institute for Materials Science (NIMS), and from Research Institute for Information Technology at Kyushu University under the category of General Projects on the supercomputer GENKAI. M.C. acknowledges GENCI for providing computational resources on the CEA-TGCC Irene supercomputer cluster under project number A0170906493 and the TGCC special session. M. C. is grateful to EuroHPC for the computational grant EHPC-EXT-2024E01-064 allocated on Leonardo booster partition.
K.N. acknowledges financial support from MEXT Leading Initiative for Excellent Young Researchers (Grant No. JPMXS0320220025), from Murata Science and Education Foundation (Grant No. M24AN006), and from Japan Science and Technology Agency (JST), PRESTO (Grant No. JPMJPR24J9). M. C. thanks the European High Performance Computing Joint Undertaking (JU) for the partial support through the "EU-Japan Alliance in HPC" HANAMI project (Hpc AlliaNce for Applications and supercoMputing Innovation: the Europe - Japan collaboration).
Associated Content
Data Availability Statement
The ab initio QMC package used in this work, TurboRVB, is available from its GitHub repository https://github.com/sissaschool/turborvb.
All data necessary for reproducing the results reported in this work is available in the following GitHub repository https://github.com/giacomotenti/unbiased_qmc_ML
Appendix A Appendix
A.1 Kernel regression models parameters
In this Appendix we provide further details on the hyperparameters and specific kernels used for the models described in Sec. 3.2. The KRR models use a kernel as the one described in Ref. 19. Each environment is described by a density function written as a sum of Gaussian functions (with spread $\epsilon$) multiplied by a localized function $f_{c}$ depending on a cutoff radius $r_{c}$, i.e.,
$\rho\left(\mathbf{r},\mathcal{R}_{i}\right)\propto\sum_{\left|\mathbf{R}_{ij}\right|\leq r_{c}}f_{c}\left(\left|\mathbf{R}_{ij}\right|\right)\exp\left(-\frac{\left|\mathbf{r}-\mathbf{R}_{ij}\right|^{2}}{\epsilon}\right).$ (13)
A similarity kernel between $\mathcal{R}_{i}$ and $\mathcal{R}_{j}$ is then defined by averaging the overlap between the two densities over a discrete set of $N_{\textrm{sym}}$ symmetry operations $\hat{U}_{k}$:
$\mathcal{K}\left(\mathcal{R}_{i},\mathcal{R}_{j}\right)=\frac{1}{N_{\textrm{sym}}}\sum_{k=1}^{N_{\mathrm{sym}}}\left[\int d^{3}\mathbf{r}\rho(\mathbf{r};\mathcal{R}_{i})\rho(\mathbf{r};\hat{U}_{k}\mathcal{R}_{j})\right]^{n}.$ (14)
Finally, the normalized kernel is written as
$K\left(\mathcal{R}_{i},\mathcal{R}_{j}\right)=\left[\frac{\mathcal{K}\left(\mathcal{R}_{i},\mathcal{R}_{j}\right)}{\mathcal{K}\left(\mathcal{R}_{i},\mathcal{R}_{i}\right)\mathcal{K}\left(\mathcal{R}_{j},\mathcal{R}_{j}\right)}\right]^{\eta}.$ (15)
In the application presented here, we chose $r_{c}=4$ Bohr and $\epsilon=1.5$ Bohr2 in Eq. (13). In Eq. (14) we averaged over the cubic symmetry group (here appropriate since we are using cubic cells) and took $n=2$. In Eq. (15) a value of $\eta=2$ was employed.
For the local energy (Eq. (4)) we also added an additional pair-wise term as done in Ref. 20, using cubic spline functions 58 defined on $10$ equally spaced grid points between $0.3$ Bohr and $5.5$ Bohr.
Finally, the furthest point sampling method was used to select $N_{\textrm{env}}=6000$ local environments among the training configurations.
A.2 Cutoff in the correction error
[…]
To manage these erratic components, in this study, we retained the biased force values in the corrected data set whenever three standard deviations exceeded a given threshold, set here to $0.015$ Ha/Bohr. As shown in Fig. 6, this threshold corresponds to a minimum in the standard deviation distribution for the corrected forces, indicating the onset of departure from normality. A similar strategy was employed for pressure, using a threshold of $1.5\times 10^{-5}$ a.u. We confirmed that forces and pressures with error bars below the threshold are unbiased, consistent with their corresponding potential energy surfaces (PESs), as illustrated in Figs.1 and 2. This choice was further validated by analyzing the behavior of $\Delta_{E}$, $\Delta_{F}$, and $\Delta_{P}$ (see Eq.(12)) as a function of the relative weight $W_{F}/W_{E}$ in the loss function, using models trained on datasets corrected with different thresholds for $\sigma_{F}$. The results, depicted in Fig. 7, confirm that the threshold of $0.015$ Ha/Bohr yields the best overall performance. Indeed, although $\Delta_{E}$ and $\Delta_{P}$ both steadily decrease with increasing threshold values (for fixed $W_{F}/W_{E}$), the accuracy of forces deteriorates significantly when corrections with large error bars (corresponding to the largest threshold, i.e. $3\sigma_{F}\leq 0.04$ Ha/Bohr) are included.
Figure 6: Distribution of three standard deviations for the corrected forces. The inset zooms-in its tail, to highlight the outliers. The vertical line indicates the $3\sigma$ threshold chosen in our application, i.e., $0.015$ Ha/Bohr.
Figure 7: $\Delta_{X}$ for energy, forces, and pressure, as a function of the force/energy weight ratio in the loss function used for training (expressed in atomic units), for models trained on different versions of the corrected dataset. Each dataset corresponds to a different value of the $3\sigma_{F}$ threshold used to accept/discard the correction. The thresholds in the legend are reported in Ha/Bohr. For pressure, a threshold of $1.5\times 10^{-5}$ a.u. is used. The weight on the pressure was set to zero for all models.
A.3 MACE models comparison
Here, we further support the conclusions made in Sec. 3.2 by carrying out the same analysis using MACE 35. This is a completely different MLIP architecture, because it is a framework based on equivariant message-passing neural networks with high-body order messages. The scope of this complementary analysis is to demonstrate that the behavior observed with the KRR models is general and only due to the training set consistency, and not linked to any particular implementation of the MLIP. The models consider $256$ inner invariant features and a cutoff of $3$ (effectively doubled when considering the message passing step). For each dataset, both training and test sets are the same as the one considered before for the KRR models in Sec. 3.2. The Adam optimizer 60 is used to find the models parameters. In particular, at each step the gradient of the loss function is evaluated on a subset of the training set ("batch") of $N_{b}=4$ configurations. The results for the different datasets are reported in Fig. 8, where we show the behavior of $\Delta_{X}$, defined in Eq. (12), for energy, forces, and pressure.
Figure 8: Relative RMSE variation $\Delta_{X}$ (Eq. 12) for energy, forces, and pressure, as a function of the force/energy weight ratio in the loss function (expressed in atomic units) for MACE MLIPs trained on the biased, corrected and reference datasets, respectively. The weight on the pressure is set to zero for all models.
They are consistent with those previously discussed in Sec. 3.2, showing the effect of the SCE and the importance of correcting it regardless of the particular choice for the MLIP framework.
A.4 Molecular dynamics simulation details
In this Appendix, we give further details on the MD simulations discussed in Sec. 3.3.
At each step, the energy, forces, and pressure were calculated at the DFT level using the Quantum Espresso package in its GPU accelerated version 61, 62, 63 with the PBE functional, and then summed with those predicted by the different MLIPs. For the DFT simulations, a $60$ Ry plane-wave cutoff with a projector augmented wave pseudopotential 64 was used together with a $4\times 4\times 4$ Monkhorst-Pack $\mathbf{k}$-point grid. For the MD simulations we used a time step of $0.25$ fs and a Langevin thermostat 65, 41 with damping $\gamma=0.13$ fs-1.
After equilibration, we ran each simulation for about $4$ ps to obtain the equilibrium quantities. The behavior of pressure as a function of the simulation time for different densities and temperatures and for each MLIP is shown in Fig. 9.
Figure 9: Pressure as a function of simulation time for a system of $128$ hydrogen atoms at different thermodynamic conditions: (a) $r_{s}=2.16$ and temperature $T=4000K$, (b) $r_{s}=2.02$ and temperature $T=6000K$, and (c) $r_{s}=1.92$ and temperature $T=8000K$. Each solid line corresponds to an MLIP trained on the different datasets while the dashed lines are the average values.